can someone tell me if this problem is right
simplify
(4+i)(6-3i)= 24-12i+6i-3i^2
= 24-12i+6i-3(-1)
= 24-6i+3
= 27-6i

Answer by funmath(2873)

simplify
(4+i)(6-3i)= 24-12i+6i-3i^2
= 24-12i+6i-3(-1)
= 24-6i+3
= 27-6i
You're right. Keep up the good work.
Happy Calculating!!!!